The Lee Corporation uses a job-order costing system. The following data were recorded for June:

June 1 Added During June
Job Number Work in Process Inventory Direct Materials Direct Labor
235 $3,150 $860 $595
236 $2,210 $1,260 $1,435
237 $1,585 $1,720 $2,530
238 $1,125 $2,410 $3,290


Overhead is charged to production at 90% of direct materials cost. Jobs 235, 237, and 238 were completed during June and transferred to finished goods. Jobs 235 and 238 have been delivered to customers. Actual manufacturing overhead was equal to manufacturing overhead applied. There were no Finished Goods inventories at the beginning of June.

Lee's Work in Process inventory balance on June 30 was:

rev: 10_24_2017_QC_CS-103206

$6,039

$11,549

$4,905

$6,199
Answer #1

Only job 236 is in work in process at the end of the month.

Overhead applied during the month = 1,260*90% = 1,134

Total work in process= 2,210 + 1,260 + 1,435 + 1,134 = 6,039

Answer is 6,039